3. Iron can be produced by the thermite reaction according to the following equation:

Fe2O3 (s) +2Al (s) produces 2 Fe (s) + Al2O3 (s)

a) If 100.0 g of Fe2O3 reacts with 30.0 g of Al, which one will be used up first?

b) How much iron (in g) could be produced?

c) If the only result in 45.0 g what is the percent yield?
Answer #1
  • Molar mass of Fe2O3 = 159.7 g/mol
  • Molar mass of Al = 27 g/mol
  • From the given equation,

159.7 g Fe2O3 reacts with 54 g Al.

So 100.0 g Fe2O3 reacts with (54×100.0)÷159.7 g = 33.8 g Al.

  • But we have 30.0 g of Al only. Thus Al acts as limiting reagent.
  • So, Al will be used up first.

54 g Al gives 2×55.85 g iron.

So 30.0 g Al gives (2×55.85×30.0)÷54 g = 62.1 g

  • Amount of iron produced = 62.1 g.
  • ​​​​​​​% yield = (amount formed×100)÷amount predicted = (45.0×100)÷62.1 = 72.5%.